# Generating Attribute-Aware Human Motions from Textual Prompt

### v1

@@ 5 Experiments
TM2T [26] 0.479±.0.002 0.686±.0.002 0.800±.0.002 1.491±.0.033 3.787±.0.014 18.5981±.0.142 2.841±.0.085
T2M [24] 0.592±.0.002 0.778±.0.002 0.859±.0.001 1.909±.0.031 3.827±.0.018 18.856±.0.217 2.627±.0.133
MotionDiffuse [123] 0.670±.0.002 0.857±.0.002 0.928±.0.001 0.416±.0.011 2.704±.0.015 18.968±.0.150 2.435±.0.119
MDM [93] 0.330±.0.006 0.479±.0.009 0.565±.0.007 1.581±.0.361 9.482±.0.105 16.294±.0.296 4.853±.0.084
T2M-GPT [121] 0.661±.0.002 0.824±.0.002 0.889±.0.002 0.279±.0.017 3.203±.0.022 19.184±.0.190 3.170±.0.162
MLD [13] 0.660±.0.002 0.850±.0.001 0.922±.0.001 0.349±.0.011 2.673±.0.013 19.576±.0.239 1.938±.0.086
GenMoStyle [28] 0.680±.0.001 0.861±.0.002 0.925±.0.001 0.332±.0.007 2.649±.0.012 19.118±.0.206 1.588±.0.074
MoMask [27] 0.685±.0.003 0.864±.0.002 0.925±.0.001 0.245±.0.009 2.602±.0.009 18.981±.0.132 1.438±.0.084
@@
Attribute Group Method R-Precision $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\rightarrow$
@@
Figure 5: Visualization of generated motions of MoMask and AttrMoGen, which demonstrates AttrMoGen’s effectiveness in producing realistic attribute-aware motions. As shown, subjects of different ages exhibit variations in the extent and patterns of their movements.
@@
A straightforward way to incorporate human attributes is to directly supplement attribute information into the text prompts and then employ conventional text-to-motion frameworks. Take the original MoMask model as an example, we conduct two key ablations: (1) human attributes are added to text prompts during only test time but not during training (‘w/ attr test’). (2) human attributes are added to text prompts during both training and testing (‘w/ attr train’). The results are illustrated in Table LABEL:tab:ablation. As shown, directly providing attributes through text prompts during inference fails to generate high-quality motions. Including attribute information in the training texts enhances MoMask’s performance, but remains less effective than our proposed attribute-control scheme.
In addition, we analyze how each component impacts the performance of our proposed model, as well as the influence of the hyperparameters $\alpha,\lambda$ in Equation 16. For clarity, AttrMoGen without $\mathcal{L}_{entropy}$ and $\mathcal{L}_{bottleneck}$ are denoted as ‘w/o entropy’ and ‘w/o bottleneck’ in the table respectively. The results in Table LABEL:tab:ablation show that each component contributes to the overall performance.
@@
A number of representative text-to-motion methods are evaluated on the HumanAttr dataset, including approaches based on VAE [24], autoregressive models [26, 121], masked transformers [27], and diffusion [123, 93, 13]. Additionally, we adopt a style-controlled motion generation model GenMoStyle [28] as a baseline by substituting style labels with attribute labels and using text prompts to generate content codes during inference. All experiments are repeated 20 times, and the mean values along with a 95% confidence interval are reported. The overall performance is summarized in Table LABEL:tab:overall. As shown, our method outperforms competing methods in R-precision, FID, MM-Dist, and Diversity. Notably, AttrMoGen advances the FID of MoMask from $0.245$ to $0.089$ and the MM-Dist from $2.602$ to $2.266$, demonstrating that incorporating human attributes can significantly enhance the quality of generated motion. Table LABEL:tab:group further compares the performance of MoMask and AttrMoGen within each attribute group, showing that AttrMoGen consistently outperforms MoMask across all groups. Notably, the performance varies across attribute groups, reflecting the diverse motion patterns associated with each group.
To validate that AttrMoGen can effectively generate motion corresponding to user input attributes, we design an additional experiment based on attribute classification. Specifically, a human attributes classifier is pre-trained on the HumanAttr training set using ground-truth attribute labels. Then the evaluation includes two protocols: given a data entry of (motion, text, attribute), (a) motion is generated using the text input and the true attributes; (b) motion is generated using the text input and a randomly assigned attribute. Then the classifier is employed to classify the attributes of the generated motions. For an ideal model, the classification results of the generated motion should match the attribute control inputs. The results in Table 5 shows that the motion generated by AttrMoGen demonstrates higher classification accuracy, indicating that it successfully generates motions aligned with user-specified attributes.
Table 5: Attribute classification accuracy of generated motion controlled by true and shuffled attribute labels.
Attribute Group True Shuffled
@@
To illustrate AttrMoGen’s ability of attribute control, we visualized the generated motions with different attribute inputs for the same text inputs, alongside MoMask’s results for comparison. As shown in Figure 8, the extent and patterns of movements vary across subjects of different ages. This demonstrates AttrMoGen’s effectiveness in producing realistic attribute-aware motions. Demo videos showcasing AttrMoGen-generated motions are available in the supplementary materials.
Since the impact of gender on motion is hard to observe visually, we design a t-SNE visualization shown in Figure 6. First, motion features extracted by the VQVAE of MoMask and AttrMoGen are visualized. Unlike MoMask that showcases distinct gender-based clustering, our Decoup-VQVAE effectively removes attribute information, resulting in no observable gender clusters. Next, we use the pretrained attribute classifier as discussed before to extract features from motions generated by MoMask and AttrMoGen. As illustrated, AttrMoGen shows more pronounced gender-based clustering compared to MoMask, indicating better alignment of generated motions with the input attributes.
6 Conclusion
This work explores the novel direction of integrating human attributes into text-to-motion generation. The core design of our framework is a decoupling encoder inspired by Structural Causal Model, which separates action semantics from human attributes, enabling text-to-semantics prediction and attribute-controlled generation. We also present a text-to-motion dataset with attribute annotations called HumanAttr, which features subjects with diverse attributes. Extensive evaluations demonstrate the potential of generating realistic human motions based on text and attribute control. Future efforts could further expand the dataset scale and the number of subjects by leveraging mocap-free data, and add more diverse types of human attributes.
@@ Appendix A Additional Proof
[…]
Note that $-\log$ is a convex function. For any given $s$, let $g(x)=p(s|x)$, $f(x)=p(x)$, and $\phi(x)=-\log x$. Applying the above inequality, we have:
[…]
@@
Table 6: Results of ablation studies. $\uparrow$” denotes that higher is better. $\downarrow$” denotes that lower is better.
Methods R-Precision $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$
Top-1 Top-2 Top-3
MDM [93] 0.330±.0.006 0.479±.0.009 0.565±.0.007 1.581±.0.361 9.482±.0.105
   w/ attr test 0.276±.0.005 0.411±.0.007 0.504±.0.007 4.320±.0.251 10.384±.0.079
   w/ attr train 0.273±.0.005 0.407±.0.007 0.494±.0.006 1.368±.0.194 10.951±.0.079
T2M-GPT [121] 0.661±.0.002 0.824±.0.002 0.889±.0.002 0.279±.0.017 3.203±.0.022
   w/ attr test 0.498±.0.002 0.673±.0.003 0.758±.0.003 1.327±.0.039 5.374±.0.032
   w/ attr train 0.655±.0.003 0.820±.0.002 0.885±.0.002 0.276±.0.016 3.274±.0.018
MLD [13] 0.660±.0.002 0.850±.0.001 0.922±.0.001 0.349±.0.011 2.673±.0.013
   w/ attr test 0.434±.0.003 0.590±.0.003 0.672±.0.002 12.727±.0.154 6.895±.0.028
   w/ attr train 0.644±.0.002 0.835±.0.002 0.910±.0.002 0.323±.0.014 2.922±.0.017
MoMask [27] 0.685±.0.003 0.864±.0.002 0.925±.0.001 0.245±.0.009 2.602±.0.009
Figure 7: Statistics of weight (in kg), height (in cm) of the HumanAttr dataset.
Attribute Group Method R-Precision $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\rightarrow$
Top-1 Top-2 Top-3 Generated / Real
Weight (kg) 0-64 MoMask 0.644±.0.005 0.835±.0.005 0.912±.0.003 0.203±.0.011 2.538±.0.023 17.674±.0.122 / 18.329±.0.244
AttrMoGen 0.652±.0.005 0.850±.0.004 0.920±.0.002 0.126±.0.007 2.388±.0.014 18.062±.0.136 / 18.329±.0.244
65-79 MoMask 0.653±.0.003 0.840±.0.003 0.914±.0.003 0.184±.0.013 2.379±.0.015 17.049±.0.183 / 17.471±.0.153
AttrMoGen 0.668±.0.004 0.854±.0.003 0.925±.0.003 0.107±.0.007 2.249±.0.018 17.165±.0.200 / 17.471±.0.153
80+ MoMask 0.646±.0.007 0.837±.0.005 0.913±.0.004 0.332±.0.020 2.439±.0.024 17.620±.0.148 / 17.962±.0.184
AttrMoGen 0.646±.0.007 0.835±.0.005 0.907±.0.003 0.244±.0.011 2.437±.0.025 18.045±.0.211 / 17.962±.0.184
Height (cm) 0-169 MoMask 0.633±.0.004 0.830±.0.005 0.906±.0.004 0.212±.0.012 2.562±.0.018 18.181±.0.150 / 18.630±.0.242
AttrMoGen 0.652±.0.005 0.842±.0.003 0.913±.0.003 0.112±.0.010 2.400±.0.018 18.408±.0.218 / 18.630±.0.242
170-179 MoMask 0.652±.0.006 0.843±.0.004 0.915±.0.003 0.241±.0.012 2.406±.0.021 16.981±.0.131 / 17.269±.0.194
AttrMoGen 0.664±.0.004 0.855±.0.003 0.923±.0.003 0.166±.0.008 2.330±.0.015 17.006±.0.153 / 17.269±.0.194
180+ MoMask 0.636±.0.006 0.822±.0.005 0.897±.0.004 0.176±.0.013 2.395±.0.019 16.765±.0.257 / 16.924±.0.197
AttrMoGen 0.639±.0.005 0.828±.0.004 0.904±.0.004 0.123±.0.012 2.267±.0.023 17.040±.0.166 / 16.924±.0.197
Overall MoMask 0.666±.0.003 0.849±.0.002 0.917±.0.001 0.169±.0.008 2.465±.0.008 17.499±.0.183 / 17.992±.0.205
AttrMoGen 0.675±.0.004 0.857±.0.002 0.924±.0.002 0.116±.0.004 2.332±.0.010 17.666±.0.175 / 17.992±.0.205
Figure 8: Visualization of generated motions of MoMask and AttrMoGen, which demonstrates AttrMoGen’s effectiveness in producing realistic attribute-aware motions.
Appendix B Additional Experimental Results
@@
To investigate the effect of directly incorporating attribute information into text prompts, we conduct experiments across several baselines [93, 121, 13, 27]. In Table LABEL:tab:ablation, ‘w/ attr test’ indicates that human attributes are added to text prompts during only test time but not during training, while ‘w/ attr train’ indicates that human attributes are added to text prompts during both training and testing. The results demonstrate that including attribute information during training generally improves the FID scores for most baselines. However, this approach remains less effective compared to our proposed AttrMoGen.
@@
Also, an additional experiment is conducted on a data subset containing human weight and height annotations, following the standard evaluation protocol. For attribute control, weight (in kg) is categorized into three groups: 0-64 as group 0, 65-79 as group 1, and 80+ as group 2. Height (in cm) is similarly divided into three groups: 0-169 as group 0, 170-179 as group 1, and 180+ as group 2. Figure 7 illustrates the distribution of weight and height within the HumanAttr dataset. The results, presented in Table LABEL:tab:group, compare the overall performance of MoMask and AttrMoGen, as well as their performance within each attribute group, showing that AttrMoGen consistently outperforms MoMask.
Figure 9: Counterfactual motions visualization. The upper-left figure depicts the original motion while the remaining are counterfactuals decoded from same semantics but different attributes.
Appendix C More Visualizations

### v2

@@ Introduction
TM2T (Guo et al. 2022b) 0.479±.0.002 0.686±.0.002 0.800±.0.002 1.491±.0.033 3.787±.0.014 18.598±.0.142 2.841±.0.085
T2M (Guo et al. 2022a) 0.592±.0.002 0.778±.0.002 0.859±.0.001 1.909±.0.031 3.827±.0.018 18.856±.0.217 2.627±.0.133
MotionDiffuse (Zhang et al. 2022) 0.670±.0.002 0.857±.0.002 0.928±.0.001 0.416±.0.011 2.704±.0.015 18.968±.0.150 2.435±.0.119
MDM (Tevet et al. 2022b) 0.330±.0.006 0.479±.0.009 0.565±.0.007 1.581±.0.361 9.482±.0.105 16.294±.0.296 4.853±.0.084
T2M-GPT (Zhang et al. 2023) 0.661±.0.002 0.824±.0.002 0.889±.0.002 0.279±.0.017 3.203±.0.022 19.184±.0.190 3.170±.0.162
MLD (Chen et al. 2023) 0.660±.0.002 0.850±.0.001 0.922±.0.001 0.349±.0.011 2.673±.0.013 19.576±.0.239 1.938±.0.086
GenMoStyle (Guo et al. 2024b) 0.680±.0.001 0.861±.0.002 0.925±.0.001 0.332±.0.007 2.649±.0.012 19.118±.0.206 1.588±.0.074
MoMask (Guo et al. 2024a) 0.685±.0.003 0.864±.0.002 0.925±.0.001 0.245±.0.009 2.602±.0.009 18.981±.0.132 1.438±.0.084
@@
Experiments
Attribute Group Method R-Precision $\uparrow$ FID $\downarrow$ $M$-Dist $\downarrow$ Diversity $\rightarrow$
@@
Implementation and Evaluation Details.
For the model architecture, MoMask (Guo et al. 2024a) is adopted due to its superior performance and efficiency. To ensure a fair comparison, all experimental settings and hyperparameters strictly follow those in MoMask. The weight coefficients $\alpha,\lambda$ of Decoup-VQVAE in Equation 16 are set to $0.01,0.5$ empirically. For attribute control, we categorize the attributes into discrete labels to accommodate sub-datasets that only provide age groups (e.g., Nymeria) and enable in-depth analyses based on categorization. Gender is labeled 0 (male) or 1 (female); age is grouped as 0 (5–18), 1 (19–35), 2 (36–59), and 3 (60–88). One-hot encodings of these labels are fused via MLPs, then concatenated with semantic embeddings for decoding. For the evaluation metrics, we follow the standard protocol (Guo et al. 2022a) and report the following metrics: FID, R-Precision, Multimodal Distance (MM-Dist), Diversity and Multimodality.
Figure 5: Visualization of generated motions of MoMask and AttrMoGen. As shown, subjects of different attributes exhibit variations in the extent and patterns of movements.
@@
A straightforward way to incorporate human attributes is to directly supplement attribute information into the text prompts and then employ conventional text-to-motion frameworks. Take the original MoMask model as an example, we conduct two key ablations: (1) human attributes are added to text prompts during only test time but not during training (‘w/ attr test’). (2) human attributes are added to text prompts during both training and testing (‘w/ attr train’). The results are illustrated in Table 2. As shown, directly providing attributes through text prompts during inference fails to generate high-quality motions. Including attribute information in the training texts enhances MoMask’s performance, but remains less effective than our proposed attribute-control scheme.
In addition, we analyze how each component impacts the performance of our proposed model, as well as the influence of the hyperparameters $\alpha,\lambda$ in Equation 16. For clarity, AttrMoGen without $\mathcal{L}_{entropy}$ and $\mathcal{L}_{bottleneck}$ are denoted as ‘w/o entropy’ and ‘w/o bottleneck’ in the table respectively. The results are presented in Table 2.
@@
A number of representative text-to-motion methods are evaluated on the HumanAttr dataset, including approaches based on VAE, autoregressive models, masked transformers, and diffusion. Additionally, we adopt a style-controlled motion generation model GenMoStyle (Guo et al. 2024b) as a baseline by substituting style labels with attribute labels and using text prompts to generate content codes during inference. All experiments are repeated 20 times, and the mean values along with a 95% confidence interval are reported. The overall performance is summarized in Table 3. As shown, our method outperforms competing methods in R-precision, FID, MM-Dist, and Diversity. Notably, AttrMoGen advances the FID of MoMask from $0.245$ to $0.089$ and the MM-Dist from $2.602$ to $2.266$, demonstrating that incorporating human attributes can significantly enhance the quality of generated motion. Table 4 further compares the performance of MoMask and AttrMoGen within each attribute group, showing that AttrMoGen consistently outperforms MoMask across all groups.
The effectiveness of attribute control is validated as follows. First, we pretrain an attribute classifier on the training set. We then use this classifier to classify the attributes of generated motions. For an ideal generation model, the predicted attributes should match the input controls. We adopt two testing protocols: (a) generating motion from (text, attribute) pairs drawn from the original test set, to evaluate performance on the original data distribution; and (b) generating motion from text with randomly assigned attributes, to assess generalization beyond the original dataset. Table 5 shows that AttrMoGen demonstrates higher classification accuracy, indicating that it successfully generates motions aligned with user-specified attributes.
Attribute Group True Random
@@
Table 5: Attribute classification accuracy of generated motion controlled by true and random attribute labels.
@@
To illustrate AttrMoGen’s ability of attribute control, we visualized the generated motions with different attribute inputs for the same text inputs, alongside MoMask’s results for comparison. As shown in Figure 9, the extent and patterns of movements vary across different attributes, which validate AttrMoGen’s effectiveness in producing attribute-aware motions.
Since the impact of gender on motion is hard to observe visually, we design a t-SNE visualization in Figure 6. First, motion features extracted by the VQVAE of MoMask and AttrMoGen are visualized. Unlike MoMask that showcases distinct gender-based clustering, Decoup-VQVAE removes attribute cues, resulting in no observable gender clusters. Then we feed MoMask‑ and AttrMoGen‑generated motions into a pretrained feature extractor and visualize the features. Here, AttrMoGen exhibits stronger gender‑based clustering than MoMask, indicating better alignment of generated motions with the input attributes.
Conclusion
This work explores the novel direction of integrating human attributes into text-to-motion generation. The core design of our framework is a decoupling encoder inspired by Structural Causal Model, which separates action semantics from human attributes, enabling text-to-semantics prediction and attribute-controlled generation. We also present a text-to-motion dataset with attribute annotations called HumanAttr, which features subjects with diverse attributes. Extensive evaluations demonstrate the potential of generating realistic human motions based on text and attribute control.
Acknowledgements
The research is supported by a grant from Bytedance (No.CT20250811106734).
@@ Appendix A Additional Proof
By Jensen’s inequality, for any real-valued convex function $\phi$ and any probability density function $f$ satisfying $\int_{x}f(x)dx=1$, the following inequality holds:
$\phi(\int_{x}g(x)f(x)dx)\leq\int_{x}\phi(g(x))f(x)dx.$ (18)
Note that $\log$ is a convex function. For any given $s$, let $g(x)=p(s|x)$, $f(x)=p(x)$, and $\phi(x)=-\log x$. Applying the above inequality, we have:
$\log(\int_{x}p(s|x)p(x)dx)\leq\int_{x}-\log(p(s|x))p(x)dx,\forall s,$ (19)
@@
Appendix B Additional Details
The dataset is split into training, validation and testing sets in an 80%, 5%, and 15% ratio. All data preprocessing protocols follow those used in HumanML3D (Guo et al. 2022a), and the input dimension for the encoder is 263. For the model architecture, MoMask (Guo et al. 2024a) is adopted due to its superior performance and efficiency. This architecture consists of a residual VQVAE, a masked transformer, and a residual transformer. The network architecture for the proxy classifier $h$ is the same as VQVAE encoder. To ensure a fair comparison, all experimental settings and hyperparameters (including seed, optimizer, epochs, scheduler, batch size, codebook size, etc.) strictly follow those in MoMask (Guo et al. 2024a). Training is performed on a single RTX 4090 GPU, requiring 8 hours for the VQVAE and 26 hours for the generative transformer.
For the evaluation metrics, we follow the standard protocol (Guo et al. 2022a) and report the following metrics: (1) Frechet Inception Distance (FID) (Heusel et al. 2017), measuring the distance between feature distributions of generated and real motions; (2) R-Precision, the accuracy for motion-to-text retrieval in a batch of 32 samples; (3) Multimodal Distance (MM-Dist), measuring the Euclidean distance between the feature vectors of text and motion; (4) Diversity, which calculates the average Euclidean distance between features of 300 randomly sampled motion pairs; (5) Multimodality, assessing the diversity of generated motions for the given text. The feature extractor used for evaluation is pre-trained on the HumanAttr dataset, following the same method as in (Guo et al. 2022a).
Regarding how action labels in certain sub-datasets are expanded into full text, there are two cases: (a) detailed descriptions of movements for each action class are provided in the database, which can be directly used (K3DA); (b) action labels themselves are already descriptive phrases; we simply add appropriate subjects to form complete sentences (ETRI, Kinder-gator).
Appendix C Additional Experimental Results
@@
To investigate the effect of directly incorporating attribute information into text prompts, we conduct experiments across several baselines (Tevet et al. 2022b; Zhang et al. 2023; Chen et al. 2023; Guo et al. 2024a). In Table 2, ‘w/ attr test’ indicates that human attributes are added to text prompts during only test time but not during training, while ‘w/ attr train’ indicates that human attributes are added to text prompts during both training and testing. The results demonstrate that including attribute information during training generally improves the FID scores for most baselines. However, this approach remains less effective compared to our proposed AttrMoGen.
@@
Also, an additional experiment is conducted on a data subset containing human weight and height annotations, following the standard evaluation protocol. For attribute control, weight (in kg) is categorized into three groups: 0-64 as group 0, 65-79 as group 1, and 80+ as group 2. Height (in cm) is similarly divided into three groups: 0-169 as group 0, 170-179 as group 1, and 180+ as group 2. Figure 7 illustrates the distribution of weight and height within the HumanAttr dataset. The results, presented in Table 4, compare the overall performance of MoMask and AttrMoGen, as well as their performance within each attribute group, showing that AttrMoGen consistently outperforms MoMask.
Appendix D More Visualizations